The Marlow Sales and Distribution Co. needs ​$540 comma 000 for the​ 3-month period ending September​ 30, 2015. The firm has explored two possible sources of credit.

a. Marlow has arranged with its bank for a ​$540 comma 000 loan secured by its accounts receivable. The bank has agreed to advance Marlow 75 percent of the value of its pledged receivables at a rate of 9 percent plus a 1 percent fee based on all receivables pledged.​ Marlow's receivables average a total of​ $1 million​ year-round.

b. An insurance company has agreed to lend the ​$540 comma 000 at a rate of 7 percent per​ annum, using a loan secured by​ Marlow's inventory of salad oil. A​ field-warehouse agreement would be​ used, which would cost Marlow ​$2 comma 000 a month. Which source of credit should Marlow​ select? Explain. Note​: Assume a​ 30-day month and​ 360-day year.
Answer #1

a. Cost of bank loan for 3 months:

loan amount*(interest rate+fee)*90/360

$540,000*(0.09+0.01)*90/360 = $540,000*0.10*90/360 = $54,000*90/360 = $4,860,000/360 = $13,500

Cost of bank loan for 3 months is $13,500.

b. Cost of loan from insurance company:

[$540,000*(0.07)*90/360] + $2,000*3 = ($37,800*90/360) + $6,000 = ($3,402,000/360) + $6,000 = $9,450 + $6,000 = $15,450

Cost of loan from insurance company for 3 months is $15,450.

So, Marlow should select loan from its bank source of credit as it's cost of $13,500 is lower than cost of loan from insurance company of $15,450.